OpenStudy (anonymous):

If Max spent 1/4 of his budget for drywall, 1/10 of his budget for concrete, 1/6 of his budget for trim, and 1/12 for hardware, what fraction of Max’s total budget has been spent? (reduce your answer to lowest terms)

OpenStudy (anonymous):

What is 36/60 reduced?

OpenStudy (anonymous):

\[\frac{12*3}{12*5}=\frac{3}{5}\]

OpenStudy (anonymous):

Did you got it ? :)

OpenStudy (anonymous):

yes. Thank you

OpenStudy (anonymous):

You are welcome.. :)